# Agent-X: Evaluating Deep Multimodal Reasoning in Vision-Centric Agentic Tasks

### v1

@@ Appendix C Metrics and Implementation Details
C.1 Error Categorization Methodology
@@
A legitimate concern is that tasks initially seeded by an LMM might advantage that same model family at evaluation time. We therefore executed a four-stage leakage-mitigation protocol:
@@
Appendix D Evaluation of newly–added models
Table 7 extends our benchmark with four open–source [^4] All systems come from publicly released weight checkpoints that can be run locally; we did not fine–tune them for Agent-X. vision LMMs (Pixtral--12B, LLaMA-3.2-11B-Vision, Kimi-VL-A3B-Thinking and mPLUG-Owl3-7B)and reports their scores under two independent evaluators: the proprietary GPT-4o [21] (“GPT judge”) and the open-source Qwen-14B [6] (“Qwen judge”). Three main takeaways emerge.
@@ Appendix D Evaluation of newly–added models
Under the Qwen rubric, Kimi-VL obtains the highest $G_{s}$, $T_{p}$, $F_{acc}$, $S_{acc}$, $G_{acc}$ and the best macro toolset-accuracy ($T^{s}_{acc}=0.62$). Even the stricter GPT judge still awards Kimi the second-best scores for grounding and semantic accuracy, confirming that Kimi-VL is presently the most well-rounded openly available model on our benchmark.
Table 7: Results on Agent–X with additional models (GPT- and Qwen-14B [7] as judges). We report performance across three evaluation modes: Step-by-Step, Deep Reasoning, and Outcome. Metrics: $\mathbf{G_{s}}$ (Grounding Score), $\mathbf{T_{p}}$ (Tool Precision), $\mathbf{T_{acc}}$ (Tool Accuracy), $\mathbf{F_{acc}}$ (Faithfulness Accuracy), $\mathbf{C_{s}}$ (Context Score), $\mathbf{F_{p}}$ (Factual Precision), $\mathbf{S_{acc}}$ (Semantic Accuracy), $\mathbf{G_{acc}}$ (Goal Accuracy), $\mathbf{G_{a}^{*}}$ (Goal Accuracy/ImgGen), and $\mathbf{T^{s}_{acc}}$ (Toolset Accuracy). The best value in each column is in bold, the second best is underlined.
Model Step-by-Step Deep Reasoning Outcome
GPT as a judge
Pixtral-12B [3] 0.12 0.20 0.63 0.45 0.19 0.26 0.34 0.07 0.55 0.54
LLaMA-3.2-11B-Vision [17] 0.03 0.15 0.14 0.70 0.08 0.70 0.24 0.07 0.26 0.42
Kimi-VL-A3B-Thinking [58] 0.26 0.19 0.59 0.62 0.42 0.52 0.65 0.29 0.29 0.48
mPLUG-Owl3-7B-240728 [73] 0.10 0.14 0.30 0.49 0.25 0.32 0.37 0.11 0.26 0.50
Qwen-14B as a judge
Pixtral-12B [3] 0.30 0.17 0.61 0.59 0.50 0.42 0.58 0.10 0.68 0.58
LLaMA-3.2-11B-Vision [17] 0.16 0.06 0.12 0.49 0.17 0.74 0.20 0.10 0.11 0.15
Kimi-VL-A3B-Thinking [58] 0.47 0.20 0.59 0.79 0.64 0.68 0.74 0.35 0.60 0.62
mPLUG-Owl3-7B-240728 [73] 0.30 0.11 0.31 0.59 0.48 0.48 0.56 0.16 0.45 0.48
@@ Appendix E Cross-Judge Consistency Study
Table 8: Additional results on Agent-X (Qwen-14B [6] as judge). We report performance across three evaluation modes: Step-by-Step, Deep Reasoning, and Outcome. Metrics: $\mathbf{G_{s}}$ (Grounding Score), $\mathbf{T_{p}}$ (Tool Precision), $\mathbf{T_{acc}}$ (Tool Accuracy), $\mathbf{F_{acc}}$ (Faithfulness Accuracy), $\mathbf{C_{s}}$ (Context Score), $\mathbf{F_{p}}$ (Factual Precision), $\mathbf{S_{acc}}$ (Semantic Accuracy), $\mathbf{G_{acc}}$ (Goal Accuracy), $\mathbf{G_{a}^{*}}$ (Goal Accuracy/ImgGen), and $\mathbf{T^{s}_{acc}}$ (Toolset Accuracy). The best value in each column is in bold, the second best is underlined.
Model Step-by-Step Deep Reasoning Outcome
Open-source
Phi-4-VL-Instruct [1] 0.27 0.11 0.32 0.54 0.39 0.59 0.46 0.16 0.35 0.39
InternVL2.5-8B [12] 0.38 0.16 0.49 0.63 0.51 0.61 0.55 0.29 0.53 0.53
Gemma-3-4B [57] 0.50 0.24 0.67 0.74 0.66 0.59 0.74 0.30 0.68 0.68
InternVL3-8B [81] 0.41 0.16 0.51 0.71 0.61 0.60 0.69 0.23 0.51 0.62
VideoLLaMA3-7B [75] 0.39 0.15 0.40 0.68 0.56 0.60 0.68 0.27 0.53 0.56
Qwen2.5-VL-7B [8] 0.51 0.27 0.63 0.77 0.66 0.64 0.77 0.37 0.62 0.67
Closed-source
Gemini-1.5-Pro [56] 0.57 0.36 0.80 0.82 0.73 0.76 0.63 0.05 0.77 0.71
Gemini-2.5-Pro [55] 0.63 0.40 0.84 0.86 0.76 0.80 0.83 0.50 0.74 0.72
GPT-4o [21] 0.46 0.27 0.63 0.72 0.59 0.75 0.69 0.44 0.48 0.56
OpenAI-o4-mini [45] 0.63 0.35 0.86 0.89 0.78 0.79 0.88 0.53 0.64 0.69
@@
Agreement on the global winner: Both judges agree that Gemini-2.5-Pro is the overall champion. When scored by the open-source Qwen-14B judge it achieves the highest values for $T_{acc}$, $F_{p}$, $S_{acc}$ and $T^{s}_{acc}$; the GPT-4o rubric yields the same outcome, with Gemini-2.5-Pro either retaining first place or falling only marginally to second. This convergence supports our claim that Gemini-2.5-Pro currently sets the bar for vision-centric agents on Agent-X. OpenAI systems show identical trends: Across both evaluations OpenAI-o4-mini secures the top $T_{acc}$ score (0.86 vs. 0.89), while GPT-4o leads in $G_{s}$ and $F_{p}$. Although small numerical variations arise from different random seeds inside each LMM judge, the relative ordering is preserved, confirming the stable performance gap between the two OpenAI releases [21, 2, 45]. Open-source models follow the same pattern: Community systems display similar consistency: Qwen2.5-VL-7B is ranked first or second by both judges for $G_{s}$, $T_{p}$, $F_{acc}$ and $T^{s}_{acc}$, whereas InternVL3-8B systematically outperforms VideoLLaMA3-7B. Absolute scores differ slightly, but no pairwise ranking is ever inverted.
@@
VideoLLaMA3-7B [75] 0.44 0.40 0.66 0.68 0.55 0.67 0.68 0.52 0.64 0.54
InternVL3-8B [81] 0.50 0.50 0.68 0.71 0.66 0.70 0.68 0.58 0.83 0.63
Qwen2.5-VL-7B [6] 0.59 0.62 0.82 0.76 0.67 0.78 0.77 0.68 0.85 0.70
@@
Table 9 reports scores assigned by two human graders who independently inspected all 120 test episodes, then reached consensus after discussion. [^5] The annotators followed the same rubric used by the automatic judges, but were allowed to replay tool outputs and visual inputs for verification; see Appendix § G. Three observations stand out.
@@
    Macro–level agreement with automatic judges. The human ranking exactly reproduces the pattern seen with Qwen-14B and GPT-4o evaluators: Gemini-2.5-Pro $>$ Qwen2.5-VL-7B $>$ InternVL3-8B $>$ VideoLLaMA3-7B. Likewise, humans confirm the sizeable lead of Gemini-2.5-Pro on every Step-by-Step metric and on the global Outcome columns ($G_{acc}$ and $T^{s}_{acc}$).
@@
Error Type Pixtral-12B [3] VideoLLaMA3-7B [75] Qwen2.5-VL-7B [6] mPLUG-Owl3 [73]
@@
Error Type GPT-4o [21] Gemini-1.5-Pro [56] OpenAI-o4-mini [45] Gemini-2.5-Pro [55]
@@ F.2 Qualitative Analysis
Figures 7 and 6 reveal two distinct error profiles. Figure 7 shows that Qwen2.5-VL-7B often hallucinates tools (e.g., a fake ObjectCounter) and then builds confident but unfounded explanations on top of those hallucinated outputs. Repeated use of the generic SceneDescriber across video frames further erodes temporal grounding. Figure 6 highlights InternVL3-8B. While it adheres to the required step-tool format, it miscounts objects, redundantly reinvokes tools, and contradicts its own earlier steps, symptoms of weak short-term memory rather than hallucination. Together, the examples confirm our quantitative results: current open-source VLMs struggle either with tool fidelity (Qwen) or cross-step coherence (InternVL), underscoring key directions for future work.
@@
F.3 Reasoning Chain Experiments

### v2

@@ Appendix C Metrics and Implementation Details
C.1 Source Datasets
Table 7: Datasets Used Across Agent-X Environments
Environment Datasets
Autonomous Driving Cityscapes (Cordts et al., 2016), BDD100K (Yu et al., 2020), Nighttime Driving (Wood, 2020), nuScenes (Caesar et al., 2020)
Security and Surveillance Surveillance Camera Fight Dataset (Aktı et al., 2019), Anomaly Detection (Sultani et al., 2018), Abnormal Events (Lu et al., 2013), AI City Challenge 2023 (Naphade et al., 2023), YouTube-hosted surveillance footage
Math Reasoning MathVista (Lu et al., 2023), AI2D (Kembhavi et al., 2016), Math Program Synthesis (Ling et al., 2017), Math Training (Cobbe et al., 2021)
Web Navigation Mobile Interfaces (Sunkara et al., 2022), WebArena (Zhou et al., 2023)
Sports SoccerNet (Deliege et al., 2021), Sports-1M (Karpathy et al., 2014), Ego4D (Grauman et al., 2022)
Generic Visual Reasoning COCO (Lin et al., 2014), Visual Genome (Krishna et al., 2017), GTA (Wang et al., 2024)
The environments in Agent-X are constructed from a diverse set of publicly available datasets spanning autonomous driving, security and surveillance, mathematical reasoning, web navigation, sports, and generic visual reasoning. These datasets provide the multimodal contexts (images, videos, and diagrams) from which tasks are derived, ensuring both diversity and realism. In particular, they enable the benchmark to capture real-world challenges such as object detection in complex scenes, anomaly recognition, multimodal mathematical problem solving, interactive navigation, and fine-grained visual reasoning.
Table 7 summarizes the datasets used in each environment along with their references. Each dataset was selected to contribute unique scenarios that demand perception-grounded reasoning and tool-based interaction, aligning with the agentic goals of Agent-X.
C.2 Error Categorization Methodology
@@
A legitimate concern is that tasks initially seeded by an LMM might advantage that same model family at evaluation time. We therefore executed a four-stage leakage‐mitigation protocol:
@@
Table 8: Results on Agent–X with additional models (GPT‐ and Qwen‐14B (Bai et al., 2023a) as judges). We report performance across three evaluation modes: Step-by-Step, Deep Reasoning, and Outcome. Metrics: $\mathbf{G_{s}}$ (Grounding Score), $\mathbf{T_{p}}$ (Tool Precision), $\mathbf{T_{acc}}$ (Tool Accuracy), $\mathbf{F_{acc}}$ (Faithfulness Accuracy), $\mathbf{C_{s}}$ (Context Score), $\mathbf{F_{p}}$ (Factual Precision), $\mathbf{S_{acc}}$ (Semantic Accuracy), $\mathbf{G_{acc}}$ (Goal Accuracy), $\mathbf{G_{a}^{*}}$ (Goal Accuracy/ImgGen), and $\mathbf{T^{s}_{acc}}$ (Toolset Accuracy). The best value in each column is in bold, the second best is underlined.
Model Step-by-Step Deep Reasoning Outcome
GPT as a judge
Pixtral-12B (Agrawal et al., 2024) 0.12 0.20 0.63 0.45 0.19 0.26 0.34 0.07 0.55 0.54
LLaMA-3.2-11B-Vision (Grattafiori et al., 2024) 0.03 0.15 0.14 0.70 0.08 0.70 0.24 0.07 0.26 0.42
Kimi-VL-A3B-Thinking (Team et al., 2025) 0.26 0.19 0.59 0.62 0.42 0.52 0.65 0.29 0.29 0.48
mPLUG-Owl3-7B-240728 (Ye et al., 2024) 0.10 0.14 0.30 0.49 0.25 0.32 0.37 0.11 0.26 0.50
Qwen-14B as a judge
Pixtral-12B (Agrawal et al., 2024) 0.30 0.17 0.61 0.59 0.50 0.42 0.58 0.10 0.68 0.58
LLaMA-3.2-11B-Vision (Grattafiori et al., 2024) 0.16 0.06 0.12 0.49 0.17 0.74 0.20 0.10 0.11 0.15
Kimi-VL-A3B-Thinking (Team et al., 2025) 0.47 0.20 0.59 0.79 0.64 0.68 0.74 0.35 0.60 0.62
mPLUG-Owl3-7B-240728 (Ye et al., 2024) 0.30 0.11 0.31 0.59 0.48 0.48 0.56 0.16 0.45 0.48
Appendix D Evaluation of Additional models
Table 8 extends our benchmark with four open–source [^4] All systems come from publicly released weight checkpoints that can be run locally; we did not fine–tune them for Agent-X. vision LMMs (Pixtral--12B, LLaMA-3.2-11B-Vision, Kimi-VL-A3B-Thinking and mPLUG-Owl3-7B)and reports their scores under two independent evaluators: the proprietary GPT-4o (Hurst et al., 2024) (“GPT judge”) and the open-source Qwen-14B (Bai et al., 2023b) (“Qwen judge”). Three main takeaways emerge.
@@ Appendix D Evaluation of Additional models
Under the Qwen rubric, Kimi-VL obtains the highest $G_{s}$, $T_{p}$, $F_{acc}$, $S_{acc}$, $G_{acc}$ and the best macro toolset‐accuracy ($T^{s}_{acc}=0.62$). Even the stricter GPT judge still awards Kimi the second-best scores for grounding and semantic accuracy, confirming that Kimi-VL is presently the most well-rounded openly available model on our benchmark.
@@
Table 9: Additional results on Agent-X (Qwen-14B (Bai et al., 2023b) as judge). We report performance across three evaluation modes: Step-by-Step, Deep Reasoning, and Outcome. Metrics: $\mathbf{G_{s}}$ (Grounding Score), $\mathbf{T_{p}}$ (Tool Precision), $\mathbf{T_{acc}}$ (Tool Accuracy), $\mathbf{F_{acc}}$ (Faithfulness Accuracy), $\mathbf{C_{s}}$ (Context Score), $\mathbf{F_{p}}$ (Factual Precision), $\mathbf{S_{acc}}$ (Semantic Accuracy), $\mathbf{G_{acc}}$ (Goal Accuracy), $\mathbf{G_{a}^{*}}$ (Goal Accuracy/ImgGen), and $\mathbf{T^{s}_{acc}}$ (Toolset Accuracy). The best value in each column is in bold, the second best is underlined.
Model Step-by-Step Deep Reasoning Outcome
Open-source
Phi-4-VL-Instruct (Abdin et al., 2024) 0.27 0.11 0.32 0.54 0.39 0.59 0.46 0.16 0.35 0.39
InternVL2.5-8B (Chen et al., 2024a) 0.38 0.16 0.49 0.63 0.51 0.61 0.55 0.29 0.53 0.53
Gemma-3-4B (Team et al., 2024b) 0.50 0.24 0.67 0.74 0.66 0.59 0.74 0.30 0.68 0.68
InternVL3-8B (Zhu et al., 2025) 0.41 0.16 0.51 0.71 0.61 0.60 0.69 0.23 0.51 0.62
VideoLLaMA3-7B (Zhang et al., 2025a) 0.39 0.15 0.40 0.68 0.56 0.60 0.68 0.27 0.53 0.56
Qwen2.5-VL-7B (Bai et al., 2023c) 0.51 0.27 0.63 0.77 0.66 0.64 0.77 0.37 0.62 0.67
Closed-source
Gemini-1.5-Pro (Team et al., 2024a) 0.57 0.36 0.80 0.82 0.73 0.76 0.63 0.05 0.77 0.71
Gemini-2.5-Pro (Team et al., 2023) 0.63 0.40 0.84 0.86 0.76 0.80 0.83 0.50 0.74 0.72
GPT-4o (Hurst et al., 2024) 0.46 0.27 0.63 0.72 0.59 0.75 0.69 0.44 0.48 0.56
OpenAI-o4-mini (OpenAI, 2025) 0.63 0.35 0.86 0.89 0.78 0.79 0.88 0.53 0.64 0.69
@@ Appendix E Cross-Judge Consistency Study
Agreement on the global winner: Both judges agree that Gemini-2.5-Pro is the overall champion. When scored by the open‐source Qwen-14B judge it achieves the highest values for $T_{acc}$, $F_{p}$, $S_{acc}$ and $T^{s}_{acc}$; the GPT-4o rubric yields the same outcome, with Gemini-2.5-Pro either retaining first place or falling only marginally to second. This convergence supports our claim that Gemini-2.5-Pro currently sets the bar for vision-centric agents on Agent-X. OpenAI systems show identical trends: Across both evaluations OpenAI-o4-mini secures the top $T_{acc}$ score (0.86 vs. 0.89), while GPT-4o leads in $G_{s}$ and $F_{p}$. Although small numerical variations arise from different random seeds inside each LMM judge, the relative ordering is preserved, confirming the stable performance gap between the two OpenAI releases (Hurst et al., 2024; Achiam et al., 2023; OpenAI, 2025). Open-source models follow the same pattern: Community systems display similar consistency: Qwen2.5-VL-7B is ranked first or second by both judges for $G_{s}$, $T_{p}$, $F_{acc}$ and $T^{s}_{acc}$, whereas InternVL3-8B systematically outperforms VideoLLaMA3-7B. Absolute scores differ slightly, but no pairwise ranking is ever inverted.
@@
VideoLLaMA3-7B (Zhang et al., 2025a) 0.44 0.40 0.66 0.68 0.55 0.67 0.68 0.52 0.64 0.54
InternVL3-8B (Zhu et al., 2025) 0.50 0.50 0.68 0.71 0.66 0.70 0.68 0.58 0.83 0.63
Qwen2.5-VL-7B (Bai et al., 2023b) 0.59 0.62 0.82 0.76 0.67 0.78 0.77 0.68 0.85 0.70
@@
Table 10 reports scores assigned by two human graders who independently inspected all 120 test episodes, then reached consensus after discussion. [^5] The annotators followed the same rubric used by the automatic judges, but were allowed to replay tool outputs and visual inputs for verification; see Appendix § H. Three observations stand out.
@@
    Macro–level agreement with automatic judges. The human ranking exactly reproduces the pattern seen with Qwen-14B and GPT-4o evaluators: Gemini-2.5-Pro >> Qwen2.5-VL-7B >> InternVL3-8B >> VideoLLaMA3-7B. Likewise, humans confirm the sizeable lead of Gemini-2.5-Pro on every Step-by-Step metric and on the global Outcome columns ($G_{acc}$ and $T^{s}_{acc}$).
@@
Error Type Pixtral-12B (Agrawal et al., 2024) VideoLLaMA3-7B (Zhang et al., 2025a) Qwen2.5-VL-7B (Bai et al., 2023b) mPLUG-Owl3 (Ye et al., 2024)
@@
Error Type GPT-4o (Hurst et al., 2024) Gemini-1.5-Pro (Team et al., 2024a) OpenAI-o4-mini (OpenAI, 2025) Gemini-2.5-Pro (Team et al., 2023)
@@ F.2 Qualitative Analysis
Figures 6 and 5 reveal two distinct error profiles. Figure 6 shows that Qwen2.5-VL-7B often hallucinates tools (e.g., a fake ObjectCounter) and then builds confident but unfounded explanations on top of those hallucinated outputs. Repeated use of the generic SceneDescriber across video frames further erodes temporal grounding. Figure 5 highlights InternVL3-8B. While it adheres to the required step-tool format, it miscounts objects, redundantly reinvokes tools, and contradicts its own earlier steps, symptoms of weak short-term memory rather than hallucination. Together, the examples confirm our quantitative results: current open-source VLMs struggle either with tool fidelity (Qwen) or cross-step coherence (InternVL), underscoring key directions for future work.
@@
Appendix G Additional Results and Analysis
This section presents a comprehensive analysis of the proposed evaluation pipeline across multiple large multimodal models (LMMs) and judges. We report results from multi-run variance studies, inter-judge agreement, automated evaluation outcomes, and reasoning–perception ablations.
Table 12: Run-to-run variance analysis for (a) Qwen2.5-VL-7B, (b) InternVL3-8B, and (c) Gemini-2.5-Pro.
Run / Metric $\mathbf{G_{s}}$ $\mathbf{T_{p}}$ $\mathbf{T_{acc}}$ $\mathbf{F_{acc}}$ $\mathbf{C_{s}}$ $\mathbf{F_{p}}$ $\mathbf{S_{acc}}$ $\mathbf{G_{acc}}$ $\mathbf{G_{a}^{*}}$ $\mathbf{T^{s}_{acc}}$
(a) Qwen2.5-VL-7B
Run 1 0.51 0.27 0.63 0.77 0.66 0.64 0.77 0.37 0.62 0.67
Run 2 0.48 0.25 0.66 0.77 0.65 0.65 0.77 0.38 0.61 0.66
Run 3 0.48 0.26 0.66 0.76 0.65 0.65 0.76 0.38 0.62 0.66
Variance (%) 0.020 0.0067 0.020 0.0022 0.0022 0.0022 0.0022 0.0022 0.0022 0.0022
(b) InternVL3-8B
Run 1 0.41 0.16 0.51 0.71 0.61 0.60 0.69 0.23 0.51 0.62
Run 2 0.36 0.15 0.50 0.62 0.50 0.62 0.56 0.20 0.51 0.52
Run 3 0.36 0.15 0.50 0.62 0.50 0.62 0.56 0.20 0.50 0.52
Variance (%) 0.0556 0.0022 0.0022 0.1800 0.2689 0.0089 0.3756 0.0200 0.0022 0.2222
(c) Gemini-2.5-Pro
Run 1 0.63 0.40 0.84 0.86 0.76 0.80 0.83 0.50 0.74 0.72
Run 2 0.62 0.38 0.86 0.85 0.75 0.82 0.84 0.49 0.73 0.71
Run 3 0.61 0.37 0.86 0.85 0.75 0.82 0.84 0.50 0.73 0.72
Variance (%) 0.0067 0.0156 0.0089 0.0022 0.0022 0.0089 0.0022 0.0022 0.0022 0.0022
G.1 Run-to-Run Variance Analysis
To ensure the reliability of our automated evaluation setup, we perform a run-to-run consistency check on three representative models: Qwen2.5-VL-7B, InternVL3-8B, and Gemini-2.5-Pro. Table 12 summarizes the variance across three independent evaluation runs for each model. Overall, the variance remains minimal across all metrics, confirming the stability of our pipeline. For instance, Qwen2.5-VL-7B and Gemini-2.5-Pro both exhibit less than 0.02% deviation across metrics, whereas InternVL3-8B shows slightly higher fluctuation (up to 0.38%) in semantic accuracy ($S_{acc}$). A high degree of determinism is therefore ensured across runs.
Table 13: Inter-judge agreement across GPT-4o, Qwen-14B, and human judges.
Statistic / Metric $\mathbf{G_{s}}$ $\mathbf{T_{acc}}$ $\mathbf{F_{acc}}$ $\mathbf{S_{acc}}$ $\mathbf{G_{acc}}$
Pearson ($r$) 0.34 0.96 0.32 0.72 0.98
Spearman ($\rho$) -0.07 0.99 0.39 0.63 0.96
Mean Dev. 0.13 0.05 0.09 0.11 0.04
Cohen’s ($\kappa$) 0.62 0.83 0.09 0.31 1.00
A consolidated summary in Table 14 further highlights that Gemini-2.5-Pro and Qwen2.5-VL achieve near-identical results across trials, confirming that the observed differences occur only beyond the third decimal place. InternVL3-8B exhibits minor instability, attributed to its stochastic sampling process during decoding, but the overall ranking remains consistent across repetitions.
G.2 Inter-Judge Agreement
To assess the robustness and reliability of our evaluation framework, we compare the judgments produced by three distinct evaluators: GPT-4o, Qwen-14B, and human experts. Table 13 presents four complementary agreement statistics, Pearson correlation ($r$), Spearman rank correlation ($\rho$), mean deviation, and Cohen’s $\kappa$, computed across five representative metrics ($G_{s}$, $T_{acc}$, $F_{acc}$, $S_{acc}$, and $G_{acc}$).
High Alignment with Human Judgments.
The results reveal remarkably strong consistency between automated judges and human evaluators, particularly in task accuracy ($T_{acc}$) and global accuracy ($G_{acc}$). Both dimensions exhibit near-perfect correlations ($r=0.96$, $\rho=0.99$ for $T_{acc}$ and $r=0.98$, $\kappa=1.00$ for $G_{acc}$), suggesting that Qwen-14B and GPT-4o are capable of replicating human evaluative behavior with high fidelity. These metrics primarily assess correctness and goal achievement, objectively quantifiable properties that leave little room for subjective interpretation. Hence, the high agreement indicates that automated judges can reliably approximate human consensus for structured, outcome-based criteria.
Moderate Agreement in Semantic and Functional Dimensions.
For semantic accuracy ($S_{acc}$) and functional accuracy ($F_{acc}$), the correlations remain moderately high (r=0.32r=0.320.720.72), reflecting partial but imperfect alignment. These metrics rely on nuanced linguistic and reasoning cues, such as whether model responses maintain contextual consistency or functional correctness, where even expert human raters may differ in interpretation. The observed variance thus highlights the intrinsic ambiguity of multimodal reasoning evaluation rather than a deficiency in the automated judges.
Lower Agreement in Goal Success.
The goal success metric ($G_{s}$) demonstrates lower correlation values ($r=0.34$, $\rho=-0.07$), revealing that high-level subjective goals are more difficult to standardize across evaluators. This finding underscores the challenge of quantifying “success” in open-ended, multimodal tasks, where different judges may focus on distinct aspects such as reasoning depth, visual grounding, or task completion. It also motivates the need for multi-perspective evaluation frameworks that combine quantitative accuracy with qualitative interpretability.
Key Insights.
Overall, the results establish that automated evaluation using Qwen-14B achieves near-human reliability across most quantitative dimensions. The strongest alignment occurs for metrics grounded in explicit correctness (e.g., $T_{acc}$ and $G_{acc}$), demonstrating that LLM-based judges are highly effective at assessing factual and structural validity. However, metrics involving contextual, semantic, or open-ended reasoning still benefit from human oversight. This reinforces the complementary nature of human–AI evaluation synergy: while automated judges provide scalability and consistency, expert human review remains crucial for capturing subtle interpretive aspects of reasoning quality.
G.3 Automated Evaluation across Models
We benchmark a wide range of vision–language models using our automated judge, Qwen-14B. As shown in Table 15, GPT-4o achieves the highest overall performance, with strong results in task accuracy ($T_{acc}=0.45$) and functional accuracy ($F_{acc}=0.34$), followed by Qwen2.5-7B-Instruct and InternVL3-8B. The baseline Qwen1.5-7B-Chat performs significantly lower across all metrics, highlighting the benefits of multimodal fine-tuning. Interestingly, Qwen2.5-VL performs comparably on low-level metrics such as $C_{s}$ and $F_{p}$ but lags in high-level goal accuracy ($G_{acc}$). This reinforces that pure visual grounding alone is insufficient for complex multimodal reasoning.
Table 14: Summary of multi-run variance analysis using Qwen-14B as judge.
Model Max Variance (%) Observation
Qwen2.5-VL 0.02 Stable across all metrics; differences appear only in the 3rd decimal place.
InternVL3-8B 0.38 Slightly higher fluctuation in semantic accuracy, but consistent overall ranking.
Gemini-2.5-Pro 0.02 Practically identical scores across runs, confirming evaluation determinism.
G.4 Reasoning vs. Perception Analysis
Table 15: Automated evaluation results across models on Agent-X.
Model $\mathbf{Inst_{align}}$ $\mathbf{T_{acc}}$ $\mathbf{Arg_{acc}}$ $\mathbf{Sum_{acc}}$ $\mathbf{F_{acc}}$ $\mathbf{C_{s}}$ $\mathbf{F_{p}}$ $\mathbf{S_{acc}}$ $\mathbf{G_{acc}}$ $\mathbf{G_{star}}$
GPT-3.5-Turbo 0.65 0.39 0.11 0.12 0.34 0.27 0.29 0.49 0.33 0.30
GPT-4o 0.72 0.45 0.15 0.12 0.34 0.31 0.36 0.51 0.36 0.35
Qwen1.5-7B-Chat 0.24 0.09 0.04 0.12 0.09 0.02 0.08 0.35 0.34 0.25
Qwen2.5-7B-Instruct 0.46 0.29 0.02 0.13 0.15 0.09 0.10 0.39 0.33 0.26
Qwen2.5-7B-VL 0.42 0.20 0.03 0.08 0.14 0.09 0.18 0.32 0.01 0.01
InternVL3-8B 0.49 0.28 0.04 0.12 0.14 0.09 0.13 0.41 0.02 0.02
To isolate and quantify the contribution of reasoning to multimodal task performance, we conduct an ablation study on a representative subset of the dataset. Specifically, we compare two configurations: (1) a perception-only model that relies solely on visual understanding without explicit reasoning steps, and (2) our proposed reasoning-augmented model that incorporates structured reasoning chains during evaluation. As shown in Table 16, the inclusion of reasoning improves the overall goal accuracy ($G_{acc}$) from 0.33 to 0.43, a relative gain of approximately 30%.
This improvement reveals that reasoning plays a crucial role in bridging low-level perception with high-level task understanding. While perception-only models can identify visual entities and basic relationships, they often fail in multi-hop inference, causal understanding, or context-dependent decision making. In contrast, models equipped with explicit reasoning chains can decompose complex agentic queries into interpretable intermediate steps, verify intermediate evidence, and better align their final predictions with the intended goals.
These findings validate our design choice of incorporating reasoning-centric evaluation protocols, as they more accurately reflect an agent’s true multimodal competence rather than surface-level visual recognition. In other words, the ability to reason explicitly, not just perceive, is what allows multimodal systems to achieve coherent, explainable, and goal-directed behavior across diverse scenarios.
Table 16: Reasoning vs. Perception Ablation on Subset.
Setting $\mathbf{G_{acc}}$
Perception-only (no explicit reasoning) 0.33
Full reasoning chain (ours) 0.43
G.5 Discussion
Our analyses collectively establish three key insights:
  1. 1.
    The evaluation pipeline is deterministic and consistent (Tables 1214), ensuring reproducibility across runs.
  2. 2.
    Automated judges like Qwen-14B show strong agreement with human evaluations (Table 13), supporting scalable benchmarking.
  3. 3.
    Explicit reasoning significantly contributes to higher multimodal task performance (Table 16), emphasizing the importance of reasoning supervision in multimodal agents.
G.6 Reasoning Chain Experiments